An empty 20000 kg coal wagon is moving on a level track at 5m/s. Suddenly 5000kg of coal is dumped into it directly from above. The coal initially has zero horizontal velocity. The final speed of the wagon is

Grade:11

1 Answers

Arun
6 years ago
According to conservation of momentum-
M1 * v1 = M2* v2
20000* 5 = 25000 * v2
100000/25000 = v2
V2 = 4 m/sec